find the height of a cylinder with a surface area of 108 square meters. the radius is twice the height
Answer by mananth(16946)

Total surface area of cylinder = 2pir(r+h)
but r= 2h
SA = = 2pi*2h(2h+h)
=12pih^2
108=6pih^2
108/12 = pi h^2
108/(12*pi)=h^2

h=1.69
Check
2*pi*3.38(1.69+3.38)